# Softplus Attention with Re-weighting Boosts Length Extrapolation in Large Language Models

### v1

@@ Abstract
Large language models have achieved remarkable success in recent years, primarily due to the implementation of self-attention mechanisms. However, traditional Softmax attention suffers from numerical instability and reduced performance as the length of inference tokens increases. This paper addresses these issues by decomposing the Softmax operation into a non-linear transformation and the $l_{1}$-norm. We identify the latter as essential for maintaining model performance. By replacing the non-linear transformation with the Softplus activation function and introducing a dynamic length scale factor for different token lengths based on invariance entropy, we create a novel attention mechanism with performance better than conventional Softmax attention across various inference lengths. To further improve the length extrapolation ability of the proposed attention mechanism, we introduce a re-weighting mechanism that amplifies significant attention weights while diminishing weaker ones, enabling the model to concentrate more effectively on relevant tokens. When combined with our proposed attention mechanism, this approach demonstrates significant promise in managing longer sequences, maintaining nearly constant validation loss even at 16$\times$ the training token length while ensuring numerical stability. Our code is available at: https://github.com/iminfine/freeatten.
@@ 1 Introduction
Softmax calculates row attention scores in two steps: first, the scores are converted to a positive domain using the natural exponential function, and then each row is normalised to ensure the scores sum to one. This process inherently produces non-negative outputs, a trait long deemed vital for its success in attention mechanisms (Qin et al., 2022; Tsai et al., 2019; Katharopoulos et al., 2020; Han et al., 2023; Arora et al., 2024). However, this assumption has come under scrutiny, as studies substituting Softmax with other non-negative functions (Ramapuram et al., 2024; Li et al., 2022; Shen et al., 2023; Wortsman et al., 2023) report performance declines. Our experiments further explore this by assessing the necessity of non-negative attention scores, revealing that while non-negativity isn’t crucial, the normalisation process is key to sustaining performance.
@@
To ensure that self-attention produces distinct peaks even for very long token length, we introduce a novel re-weighting approach that enhances the model’s capacity to concentrate on relevant tokens while filtering out noise. By applying a power transformation to the attention scores and $l_{1}$-norm afterwards, this re-weighting mechanism effectively amplifies stronger attention weights and suppresses weaker ones. Empirical results demonstrate that this re-weighting mechanism can be seamlessly integrated with various attention variants, leading to improved performance. When combined with our proposed LSSA, as shown in Figure 1 it shows particular promise in managing longer sequences, maintaining nearly constant validation loss even at 16 $\times$ the training token length while ensuring numerical stability.
@@
    We propose a re-weighting approach that enhances the model’s ability to focus on relevant tokens by amplifying significant attention weights while reducing weaker ones. When combined with our proposed LSSA, this re-weighted model (LSSAR) excels particularly in length extrapolation while ensuring numerical stability.
@@ 2.3 Attention Re-weighting Mechanism
where $\mathbf{O}$ denotes a matrix where every element is equal to one  [^2] To prevent training instability due to the absence of positive values in the initial rows, the elements within first three rows of $\mathbf{O}$ were set to zero.. The mean effective value of $\mathbf{A}\otimes\mathbf{N}$ is one, leading to an expectation of effective value of $\mathbf{A}\otimes\mathbf{N}-\mathbf{O}$ being zero. The notation $\operatorname{ReLU}^{p}$ indicates that the output of the ReLU function is raised to the power of $p$, with $p$ being a positive integer. ReLU effectively masks tokens with small attention scores, while the power operation enhances the discriminability of the remaining tokens.
@@ 3 Experiments
All models were trained using a sequence length of 1024 on the fineweb-10B dataset [^3] https://huggingface.co/datasets/HuggingFaceFW/fineweb, which consists of 10.2 billion training tokens distributed across 18,865 training steps, along with 0.1 billion validation tokens.
@@ 5 Related Work
Length Extrapolation. Positional embeddings (Su et al., 2024; Chen et al., 2023; Chi et al., 2022; Kiyono et al., 2021; Golovneva et al., 2024; He et al., 2024; Huang et al., 2020; Li et al., 2023; Likhomanenko et al., 2021; Liu et al., 2023; Wang et al., 2024; Zheng et al., 2024) play a vital role in transformer architectures by providing essential information about the positions of tokens within a sequence, which is considered a key factor in enhancing length extrapolation (Kazemnejad et al., 2023). Among these embeddings, RoPE is particularly noteworthy; it forms the foundation of many modern LLMs, including GPT-4 (Achiam et al., 2023), Llama3 (Dubey et al., 2024), Qwen2 (Yang et al., 2024) and Deepseek-v3 (Liu et al., 2024a). The underlying model, GPT-2-124m, employed in all experiments in this paper, was modified to integrate RoPE, and our proposed LSSAR method significantly improves its length extrapolation capabilities. This indicates that a well-designed attention mechanism like LSSAR is essential for RoPE-based LLMs to enhance their length extrapolation, suggesting that LSSAR may serve as a general strategy to bolster the length extrapolation abilities of most contemporary LLMs.
@@ 6 Conclusion
In this paper, we present a novel attention mechanism called LSSA. The primary innovation involves decomposing the Softmax operation, used in the traditional self-attention framework, into a non-linear transformation and the $l_{1}$-norm, replacing the non-linear transformation with the Softplus activation function and introducing a dynamic length scale factor related to token length. Experimental results indicate that LSSA surpasses conventional Softmax attention across various sequence lengths. To improve the length extrapolation capabilities of LSSA, we have devised a distinctive re-weighting mechanism that, when integrated with LSSA, the resulting model LSSAR significantly enhances its ability to extrapolate over longer sequences while maintaining stability and scalability. This positions LSSAR as a compelling candidate for large-scale models. Consequently, we recommend the incorporation of LSSAR in future transformer architectures.

### v2

@@ Abstract
Large language models have achieved remarkable success in recent years, primarily due to the implementation of self-attention mechanisms. However, traditional Softmax attention suffers from numerical instability and reduced performance as the length of inference tokens increases. This paper addresses these issues by decomposing the Softmax operation into a non-linear transformation and the $l_{1}$-norm. We identify the latter as essential for maintaining model performance. By replacing the non-linear transformation with the Softplus activation function and introducing a dynamic scale factor for different token lengths based on invariance entropy, we create a novel attention mechanism with performance better than conventional Softmax attention across various inference lengths. To further improve the length extrapolation ability of the proposed attention mechanism, we introduce a fine-tuning-free re-weighting mechanism that amplifies significant attention weights while diminishing weaker ones, enabling the model to concentrate more effectively on relevant tokens without requiring retraining. When combined with our proposed attention mechanism, this approach demonstrates significant promise in managing longer sequences, maintaining nearly constant validation loss even at 16$\times$ the training token length while ensuring numerical stability. Our code is available at: https://github.com/iminfine/freeatten.
@@ 1 Introduction
Softmax calculates row attention scores in two steps: first, the scores are converted to a positive domain using the natural exponential function, and then each row is normalised to ensure the scores sum to one. This process inherently produces non-negative outputs, a trait long deemed vital for its success in attention mechanisms (Tsai et al., 2019; Katharopoulos et al., 2020; Han et al., 2023; Arora et al., 2024). However, this assumption has come under scrutiny, as studies substituting Softmax with other non-negative functions (Ramapuram et al., 2024; Li et al., 2022; Shen et al., 2023; Wortsman et al., 2023) report performance declines. Our experiments further explore this by assessing the necessity of non-negative attention scores, revealing that while non-negativity isn’t crucial, the normalisation process is key to sustaining performance.
@@
To ensure that self-attention produces distinct peaks even for very long token length, we introduce a novel fine-tuning-free re-weighting approach that enhances the model’s capacity to concentrate on relevant tokens while filtering out noise. By applying a power transformation to the attention scores and $l_{1}$-norm afterwards, this re-weighting mechanism effectively amplifies stronger attention weights and suppresses weaker ones without retraining or architectural changes. Empirical results demonstrate that this re-weighting mechanism can be seamlessly integrated with various attention variants, leading to improved performance. When combined with our proposed LSSA, as shown in Figure 1 it shows particular promise in managing longer sequences, maintaining nearly constant validation loss even at 16 $\times$ the training token length while ensuring numerical stability.
@@
    We propose a novel fine-tuning-free re-weighting approach that enhances the model’s ability to focus on relevant tokens by amplifying significant attention weights while reducing weaker ones without requiring retraining or architectural modifications. When combined with our proposed LSSA, this re-weighted model (LSSAR) excels particularly in length extrapolation while ensuring numerical stability.
@@ 2.3 Attention Re-weighting Mechanism
where $\mathbf{O}$ denotes a matrix where every element is equal to one  [^2] To prevent training instability due to the absence of positive values in the initial rows, the elements within first three rows of $\mathbf{O}$ were set to zero.. The mean effective value of $\mathbf{A}\otimes\mathbf{N}$ is one, leading to an expectation of effective value of $\mathbf{A}\otimes\mathbf{N}-\mathbf{O}$ being zero. The notation $\operatorname{ReLU}^{p}$ indicates that the output of the ReLU function is raised to the power of $p$, with $p$ being a positive integer. ReLU effectively masks tokens with small attention scores, while the power operation enhances the discriminability of the remaining tokens. The re-weighting mechanism is applied post-training, requiring no additional gradient updates, which preserves the pre-trained model’s parameters intact and avoids the computational burden of fine-tuning.
@@ 3 Experiments
All models were trained using a sequence length of 1024 on the fineweb-10B dataset (Penedo et al., 2024), which consists of 10.2 billion training tokens distributed across 18,865 training steps, along with 0.1 billion validation tokens.
@@ 5 Related Work
Length Extrapolation. Positional embeddings (Su et al., 2024; Chen et al., 2023; Chi et al., 2022; Kiyono et al., 2021; Golovneva et al., 2024; He et al., 2024; Huang et al., 2020; Li et al., 2023; Likhomanenko et al., 2021; Liu et al., 2023; Wang et al., 2024; Zheng et al., 2024) play a vital role in transformer architectures by providing essential information about the positions of tokens within a sequence, which is considered a key factor in enhancing length extrapolation (Kazemnejad et al., 2023). Among these embeddings, RoPE is particularly noteworthy; it forms the foundation of many modern LLMs, including GPT-4 (Achiam et al., 2023), Llama3 (Dubey et al., 2024), Qwen2 (Yang et al., 2024), Deepseek-v3 (Liu et al., 2024a) and DeepSeek-R1 (DeepSeek-AI et al., 2025). While approaches like Position Interpolation (Chen et al., 2023; Li et al., 2023) modify positional encodings and require costly fine-tuning on extended sequences, the proposed LSSAR achieves robust length extrapolation without retraining, directly enhancing pre-trained models through its attention mechanism. This indicates that a well-designed attention mechanism like LSSAR is essential for RoPE-based LLMs to enhance their length extrapolation, suggesting that LSSAR may serve as a general strategy to bolster the length extrapolation abilities of most contemporary LLMs.
@@ 6 Conclusion
In this paper, we present a novel attention mechanism called LSSA. The primary innovation involves decomposing the Softmax operation, used in the traditional self-attention framework, into a non-linear transformation and the $l_{1}$-norm, replacing the non-linear transformation with the Softplus activation function and introducing a dynamic length scale factor related to token length. Experimental results indicate that LSSA surpasses conventional Softmax attention across various sequence lengths. To improve the length extrapolation capabilities of LSSA, we have devised a distinctive fine-tuning-free re-weighting mechanism that, when integrated with LSSA, the resulting model LSSAR significantly enhances its ability to extrapolate over longer sequences without requiring retraining, while maintaining stability and scalability, making it a practical, cost-effective solution for enhancing long-context capabilities in pre-trained LLMs without the computational burden of fine-tuning. This positions LSSAR as a compelling candidate for large-scale models. Consequently, we recommend the incorporation of LSSAR in future transformer architectures.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.